# Cyclotomic Factors and LRS-Degeneracy

### v2

@@ 5.2 Preprocessing prior to testing LRS-degeneracy
    if $f(x)=g(x^{r})$ then $f$ is $k$-LRS-degenerate for each factor $k>1$ of $r$; if additionally $g(x)$ is $\kappa$-LRS-degenerate then $f$ is $r\kappa$-LRS-degenerate; $f$ may also be $k$-LRS-degenerate for other $k\mid r\kappa$.
@@
  • we may assume that $f$ has no cyclotomic factors: otherwise it is trivially LRS-degenerate — e.g. use the method of Section 4
@@ 5.2.1 Preprocessing: coefficient reduction
If the input is $f=a_{d}x^{d}+a_{0}$ then we can directly simplify to $x^{d}+1$. As presented, the algorithm will fail to fully simplify $f$ if it is unable to factorize $a_{0}$ or $a_{d}$.
@@
4:Compute $g\leftarrow\gcd(a_{1},a_{2},\ldots,a_{d})$triangleright $a_{0}$ is deliberately excluded
@@
12:return $f(\lambda x)$triangleright element of ${\mathbb{Z}}[x]$
@@ 5.3 Factorization and LRS-degeneracy
Let $g(x)\in{\mathbb{C}}[x]$ be not LRS-degenerate then $h(x)=g(x)\,g(-x)$ is $k$-LRS-degenerate only for $k=2$.
@@
If $g(\beta)=0$ then $g$ is $k$-LRS-degenerate. Otherwise we have $g(-\beta)=0$; so $g(x)$ has a root pair with ratio $-\alpha/\beta=-\zeta_{k}$, and hence $g(x)$ is LRS-degenerate (with order either $k/2$ or $2k$). ∎
@@
It is easy to see that if $f(x)\,g(x)$ is LRS-degenerate then so is $f(x)\,g(-x)$. Note that $f(-x)\,g(x)$ is LRS-degeneracy equivalent to $f(x)\,g(-x)$, and $f(-x)\,g(-x)$ is LRS-degeneracy equivalent to $f(x)\,g(x)$.
@@ 5.6 The modular algorithm
We now present our modular approach whose existence was hinted at in [3]. Our approach employs “sophisticated brute force”, and comprises two parts: one testing whether a polynomial is $k$-LRS-degenerate for a specific order $k$, the other choosing which orders $k$ to test. Recall that the input is content-free, square-free $f\in{\mathbb{Z}}[x]$ with $f(0){\neq}0$, degree $\geq 2$ and $\gcd(f(x),f(-x)){=}1$. We give first the algorithm then further details and justifications in the following subsections.
@@ 5.7 The modular algorithm: supporting lemma
11:              skip to next $k$ end $\triangleright$ $\bar{f}$ is canonical image of $f$ in ${\mathbb{F}}_{p}[x]$
@@ 5.8 Test for $k$-LRS-degeneracy
In the loop at Step 9 we can stop iterating at $\lfloor k/2\rfloor$ because for $j>k/2$ we have $\zeta_{k}^{j}=1/\zeta_{k}^{k-j}$, so letting $g=\gcd(\bar{f}(x),\bar{f}(\zeta_{k}^{k-j}x))$ we deduce that $\gcd(\bar{f}(x),\bar{f}(\zeta_{k}^{j}x))=u\,g(\zeta_{k}^{j}x)$ for some unit $u$. So both gcds have the same degree. Here $\bar{f}$ denotes the canonical image of $f$ in ${\mathbb{F}}_{p}[x]$.
To further reduce the chance of a false positive, we perform the analogous test for up to two more primes: the loop at Step 6 performs at most 3 iterations. If all these checks pass then “very likely” $f$ is indeed $k$-LRS-degenerate, so we return a positive result. Quantifying the probability of a false positive would require assuming a distribution on the input polynomials, but there is no obvious reasonable choice.
@@ 5.11 Timings
In Table 3 we present some timings of our implementation of Algorithm 5 (LRSDegeneracyOrders). We tried the three largest examples from [3], each of degree 24; but each example took less than 0.1s (to compute all LRS-degeneracy orders). Instead we generated some higher degree polynomials: four with uniform random integer coefficients from $-2^{10}$ to $2^{10}$ (a larger coefficient range made essentially no difference); and three products of cyclotomic polynomials which have many orders. As expected, none of the random polynomials was LRS-degenerate.
@@
$\Phi_{123}\Phi_{185}\Phi_{217}\Phi_{299}\Phi_{319}\Phi_{323}$ 1 23612361\,2361 236 3 24332433\,2433 243
In none of our tests did we observe any false positive. Also, for polynomials with larger LRS-degeneracy orders, the costs of verification dominate: for instance, we did not perform verification for the degree 1 23612361\,2361 236 polynomial as fifteen of the candidate orders are at least 22 7552275522\,75522 755, and the largest order is 103 037103037103\,037103 037.

### v3

@@ 5.2 Preprocessing prior to testing LRS-degeneracy
    if $f(x)=g(x^{r})$ then $f$ is $k$-LRS-degenerate for each factor $k>1$ of $r$; if additionally $g(x)$ is $\kappa$-LRS-degenerate then $f$ is $r\kappa$-LRS-degenerate; $f$ may also be $k$-LRS-degenerate for other factors $k\mid r\kappa$.
@@
  • we may assume that $f$ has no cyclotomic factors (with index >2>2): otherwise it is trivially LRS-degenerate — e.g. use the method of Section 4
@@ 5.2.1 Preprocessing: coefficient reduction
If the input is $f=a_{d}x^{d}+a_{0}$, with $a_{d}\,a_{0}\neq 0$, then we can directly simplify to $x^{d}+1$. As presented, the algorithm will fail to fully simplify $f$ if it is unable to factorize $a_{0}$ or $a_{d}$.
@@
4:Compute $g\leftarrow\gcd(a_{1},a_{2},\ldots,a_{d})$ $\triangleright$ $a_{0}$ is deliberately excluded
@@
12:return $f(\lambda x)$ $\triangleright$ element of ${\mathbb{Z}}[x]$
@@ 5.3 Factorization and LRS-degeneracy
Let $g(x)\in{\mathbb{C}}[x]$ be not LRS-degenerate then h(x)=g(x)g(x)h(x)=g(x)\,g(-x) is $k$-LRS-degenerate only for $k=2$.
@@
If $g(\beta)=0$ then $g$ is $k$-LRS-degenerate. Otherwise we have $g(-\beta)=0$; so $g(x)$ has a root pair with ratio $\alpha/\beta=-\zeta_{k}$, and hence $g(x)$ is LRS-degenerate (with order either $k/2$ or $2k$). ∎
@@
It is easy to see that if f(x)g(x)f(x)\,g(x) is LRS-degenerate then so is f(x)g(x)f(x)\,g(-x). Note that f(x)g(x)f(-x)\,g(x) is LRS-degeneracy equivalent to f(x)g(x)f(x)\,g(-x), and f(x)g(x)f(-x)\,g(-x) is LRS-degeneracy equivalent to f(x)g(x)f(x)\,g(x).
@@ 5.6 The modular algorithm
We now present our modular approach whose existence was hinted at in [3]. Our approach employs “sophisticated brute force”, and comprises two parts: one testing whether a polynomial is $k$-LRS-degenerate for a specific order $k$, the other choosing which orders $k$ to test. Recall that the input is content-free, square-free $f\in{\mathbb{Z}}[x]$ with $f(0){\neq}0$, degree $\geq 2$ and $\gcd(f(x),f(-x)){=}1$. The following subsections give further details and justifications for our new Algorithm 5 (LRSDegeneracyOrders).
@@ 5.8 Test for $k$-LRS-degeneracy
[…]
@@
In the loop at Step 9 we can stop iterating at $\lfloor k/2\rfloor$ because for $j>k/2$ we have $\zeta_{k}^{j}=1/\zeta_{k}^{k-j}$, so letting $g=\gcd(\bar{f}(x),\bar{f}(\zeta_{k}^{k-j}x))$ we deduce that $\gcd(\bar{f}(x),\bar{f}(\zeta_{k}^{j}x))=u\,g(\zeta_{k}^{j}x)$ for some unit $u$. So both gcds have the same degree. Here we write $\bar{f}$ to denote the canonical image of $f$ in ${\mathbb{F}}_{p}[x]$.
To further reduce the chance of a false positive, we perform the analogous test for up to two more primes: the loop at Step 6 performs at most 3 iterations. If all these checks pass then “very likely” $f$ is indeed $k$-LRS-degenerate, so we return a positive result. Quantifying the probability of a false positive would require assuming a distribution on the input polynomials, but there is no obvious, reasonable choice.
@@ 5.11 Timings
In Table 3 we present some timings of our implementation of Algorithm 5 (LRSDegeneracyOrders). We tried the three largest examples from [3], each of degree 24; but each example took less than 0.1s (to compute all LRS-degeneracy orders). Instead we generated some higher degree polynomials: four with uniform random integer coefficients from $2^{10}$ to $2^{10}$ (a larger coefficient range made essentially no difference); and three products of cyclotomic polynomials which have many orders. As expected, none of the random polynomials was LRS-degenerate.
@@
$\Phi_{123}\Phi_{185}\Phi_{217}\Phi_{299}\Phi_{319}\Phi_{323}$ 1 2361\,236 3 2433\,243
In none of our tests did we observe any false positive. Also, for polynomials with larger LRS-degeneracy orders, the costs of verification dominate: for instance, we did not perform verification for the degree 1 2361\,236 polynomial as fifteen of the candidate orders are at least 22 75522\,755, and the largest order is 103 037103\,037.